OpenStudy (anonymous):

No solution or infinite solution 6x+2y=8 2x-y=6

OpenStudy (shiraz14):

@cutegirl : 6x+2y=8 --> 3x+y=4 2x-y=6 Adding both equations, we have: 5x=10 --> x=2 Substituting x=2 into the first equation, we have: y=4-3(2) =4-6 =-2